# Anderson localization on quantum graphs coded by elements of a subshift of finite type

### v1

@@ 1. Main result
While the interiors of the intervals $I_{n,j}$ are assumed to be disjoint, we will also assume that their endponts are shared in the sense that $n$ and $n+1$ belong to all intervals $I_{n,j}$. Thus,
@@
Since $\Omega$ is a metric space, we can talk about the Borel $\sigma$-algebra of subsets of $\Omega$ and consider probability measures on $\Omega$. Let $\mu$ be a $T$-ergodic probability measure on $\Omega$. We will not work with arbitrary measures $\mu$. Instead, we will assume that $\mu$ has a bounded distorsion property. To give a formal definition of this property, we first define the cylinder sets
@@
Definition. The measure $\mu$ is said to be a measure with a bounded distorsion property, provided there is a constant $C>1$ such that
@@
Let $\Omega$ be a subshift of finite type and let $\mu$ be a $T$-ergodic measure with the bounded distorsion property. Suppose that ${\rm supp}\,(\mu)=\Omega$ and that $T$ has at least one fixed point and at least one point that is not fixed. Then for $\mu$-almost every $\omega$, the operator $H_{\omega}$ has pure point spectrum that coincides with $0,\infty)$ as a set. Moreover, there is a finite subset ${\mathfrak{X}}\subset[0,2\pi]$ that does not depend on $\omega$ such that if $E$ is an eigenvalue of $H_{\omega}$ and $\sqrt{E}-2\pi k\notin{\mathfrak{X}}$ for all $k=0,1,\dots$, then the corresponding eigenfunction decays exponentially as $|x|\to\infty.$
@@
Suppose $T:\Omega\to\Omega$ is a subshift of finite type and $\mu$ is a $T$-ergodic probability measure that has a local product structure and posseses the property ${\rm supp}\,\mu=\Omega$. Suppose $T$ has a fixed point, and at least one $\omega\in\Omega$ that is not a fixed point. Then the set ${\mathfrak{L}}(\mu)$ is finite. Moreover, if $E>0$ is such that $(\sqrt{E}-2\pi n)^{2}\notin{\mathfrak{L}}(\mu)$ for all $n=0,1,\dots$, then $L(E)>0$.
@@
has an invariant probability measure $m$ on $\Omega\times{\mathbb{R}\mathbb{P}}^{1}.$ We say that such a measure $m$ projects to $\mu$ if $m(\Delta\times{\mathbb{R}\mathbb{P}}^{1})=\mu(\Delta)$ for all Borel subsets $\Delta$ of $\Omega$. Given any $T$-invariant measure $\mu$ on $\Omega$, one can find a $(T,A^{E})$-ivariant measure $m$ that projects to $\mu$ by applying the standard Krylov-Bogolyubov trick used to construct invariant measures.
@@
Let $\mathcal{F}$ be the set of energies $E\geqslant 0$ for which there is a measure $m$ that is both a u-state and an s-state. It was shown in [27] that the set $\mathcal{F}$ is discrete. It is convenient to add the squares of integer multiples of $\pi$ to $\mathcal{F}$, and consider the union
$\tilde{\mathcal{F}}={\mathcal{F}}\cup\bigcup_{n=0}^{\infty}\{(\pi n)^{2}\}$
@@
$B_{\eta}(\tilde{\mathcal{F}})=\{E>0:\,\,{\rm dist}(E,\mathcal{F})<\eta\}.$
@@
By $T_{+}$-invariance of $\mu^{+}$, and the bounded distorsion property of the measure, for admissible $j,i_{1},\dots,i_{s}$, we have
@@ 3. Localization as a consequence of ULD
Another important relation is the one that expresses $A^{E}_{N}(\omega)$ in terms of determinants considered above. Let $u$ be the formal solution of the equation $H_{\omega}u=Eu$ satisfying the intitial conditions u(1)=0u(-1)=0 and $u(0)=1$. Then
$({\mathcal{H}}_{\omega,N}-\tilde{E})u=-\alpha_{N-1}u(N)e_{N-1}\implies 1=-\alpha_{N-1}u(N)G^{E}_{\omega,N}(0,N-1).$
@@
$u(N)=\det(\tilde{E}-{\mathcal{H}}_{\omega,N})\prod_{i=0}^{N-1}\alpha_{i}^{-1}$
Let now $v$ be the solution satisfying the intitial conditions v(1)=1v(-1)=1 and $v(0)=0$. Then $\alpha_{0}v(1)/\alpha_{-1}=1$. Therefore,
$\frac{\alpha_{0}}{\alpha_{-1}}v(N)=\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-1})\prod_{i=1}^{N-1}\alpha_{i}^{-1}.$
@@
$A^{E}_{N}(\omega)=\sqrt{\frac{\omega_{N-1}}{\omega_{-1}}}\begin{pmatrix}\det(\tilde{E}-{\mathcal{H}}_{\omega,N})\prod_{i=0}^{N-1}\alpha_{i}^{-1}&-\alpha_{-1}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-1})\prod_{i=0}^{N-1}\alpha_{i}^{-1}\\ \,&\,\\ \det(\tilde{E}-{\mathcal{H}}_{\omega,N-1})\prod_{i=0}^{N-2}\alpha_{i}^{-1}&-\alpha_{-1}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-2})\prod_{i=0}^{N-2}\alpha_{i}^{-1}\end{pmatrix}$. (3.15)
@@
Assume that $\mu$ has a bounded distorsion property. Then there is a constant $C\geqslant 1$ such that
@@
Proof of Proposition 3.7. Let $N$ and $\varepsilon$ be given. Fix $s\in{\mathbb{Z}}$ and $K\geqslant\max\{N,\ln^{2}(|s|+1)\}$. Define the set $D_{K,s}$ as the set of $\omega$ for which (3.24),(3.25) are satisfied for some $N_{1}$, $N_{2}$, $r$, and $m$ obeying the conditions $0\leqslant N_{1},N_{2}\leqslant K^{9}$, $K^{10}\leqslant r\leqslant\bar{K}$, and $m\in\{K,2K\}$. Note that
@@
Proof. For any interval $[a,b]$,
$u(n)=-G_{\omega,[a,b]}^{E}(n,a)\alpha_{a-1}u(a-1)-G^{E}_{\omega,[a,b]}(n,b)\alpha_{b}u(b+1)$ (3.44)
@@
$\begin{split}|u(\ell_{i}+s_{0})|\leqslant 2(|u(a_{i}+s_{0})|+|u(b_{i}+s_{0}-1)|)\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3})\\ \leqslant C_{u}(5+|a_{i}|+|b_{i}|)\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3})\\ \leqslant C_{u}(5+2K^{9})\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3}).\end{split}$
@@
Now we use the condition $|u(s_{0}|=1$ and apply (3.44) with $a=s_{0}+\ell_{1}+1$ and $b=s_{0}+\ell_{2}-1$ to get the estimate
$\begin{split}1=|u(s_{0})|\leqslant 2|G^{E}_{\omega,[a,b]}(s_{0},a)|\,|u(s_{0}+\ell_{1})|+2|G^{E}_{\omega,[a,b]}(s_{0},b)|\,|u(s_{0}+\ell_{2})|\\ \leqslant 2\bigl(|G^{E}_{\omega,[a,b]}(s_{0},a)|+|G^{E}_{\omega,[a,b]}(s_{0},b)|\bigr)e^{-2K^{2}}\leqslant 4\|G^{E}_{\omega,[a,b]}\|e^{-2K^{2}}.\end{split}$
@@
To move further, we need the following lemma.
@@
for all suficiently large $N_{0}$. This implies the relation

### v2

@@ 1. Main result
This paper extends findings from [27] regarding the positivity of the Lyapunov exponent in quantum graphs with a variable number of edges. By combining this positivity with large deviation estimates, we establish spectral characteristics and prove localization for these systems.
@@
While the interiors of the intervals $I_{n,j}$ are assumed to be disjoint, we will also assume that their endpoints are shared in the sense that $n$ and $n+1$ belong to all intervals $I_{n,j}$. Thus,
@@
Since $\Omega$ is a metric space, we can talk about the Borel $\sigma$-algebra of subsets of $\Omega$ and consider probability measures on $\Omega$. Let $\mu$ be a $T$-ergodic probability measure on $\Omega$. We will not work with arbitrary measures $\mu$. Instead, we will assume that $\mu$ has a bounded distortion property. To give a formal definition of this property, we first define the cylinder sets
@@
Definition. The measure $\mu$ is said to be a measure with a bounded distortion property, provided there is a constant $C>1$ such that
@@
Let $\Omega$ be a subshift of finite type and let $\mu$ be a $T$-ergodic measure with the bounded distortion property such that ${\rm supp}\,(\mu)=\Omega$. Suppose $T$ has at least one fixed point and at least one non-fixed point. Then for $\mu$-almost every $\omega$, the operator $H_{\omega}$ has pure point spectrum equal to $0,\infty)$. Furthermore, there is a finite subset ${\mathfrak{X}}\subset[0,2\pi]$, independent of $\omega$, such that for any eigenvalue $E$ of $H_{\omega}$, if $\sqrt{E}-2\pi k\notin{\mathfrak{X}}$ for all $k=0,1,\dots$, then the corresponding eigenfunction decays exponentially as $|x|\to\infty.$
@@
Let $T:\Omega\to\Omega$ be a subshift of finite type and $\mu$ be a $T$-ergodic probability measure with full support (${\rm supp}\,\mu=\Omega$) and a local product structure. Assume $T$ possesses at least one fixed point and at least one non-fixed point $\omega\in\Omega$. Then the set ${\mathfrak{L}}(\mu)$ is finite. Furthermore, if $E>0$ satisfies $(\sqrt{E}-2\pi n)^{2}\notin{\mathfrak{L}}(\mu)$ for all $n=0,1,\dots$, then $L(E)>0$.
@@
has an invariant probability measure $m$ on $\Omega\times{\mathbb{R}\mathbb{P}}^{1}.$ We say that such a measure $m$ projects to $\mu$ if $m(\Delta\times{\mathbb{R}\mathbb{P}}^{1})=\mu(\Delta)$ for all Borel subsets $\Delta$ of $\Omega$. Given any $T$-invariant measure $\mu$ on $\Omega$, one can find a $(T,A^{E})$-invariant measure $m$ that projects to $\mu$ by applying the standard Krylov-Bogolyubov trick used to construct invariant measures.
@@
Let $\mathcal{E}$ be the set of energies $E\geqslant 0$ for which there is a measure $m$ that is both a u-state and an s-state. It was shown in [27] that the set $\mathcal{E}$ is discrete. It is convenient to add the squares of integer multiples of $\pi$ to $\mathcal{F}$, and consider the union
$\tilde{\mathcal{E}}={\mathcal{E}}\cup\bigcup_{n=0}^{\infty}\{(\pi n)^{2}\}.$
@@
$B_{\eta}(\tilde{\mathcal{E}})=\{E>0:\,\,{\rm dist}(E,\tilde{\mathcal{E}})<\eta\}.$
@@
By $T_{+}$-invariance of $\mu^{+}$, and the bounded distortion property of the measure, for admissible $j,i_{1},\dots,i_{s}$, we have
@@ 3. Localization as a consequence of ULD
In the proposition below, instead of the exponent $2/3$ obtained in [2], we obtain a smaller exponent $2/7$, which is justified by our analysis of a special case.
@@
Another important relation is the one that expresses $A^{E}_{N}(\omega)$ in terms of determinants considered above. Let $u$ be the formal solution of the equation $H_{\omega}u=Eu$ satisfying the intitial conditions u(1)=0u(-1)=0 and $u(0)=1$. Define $w(n)=\sqrt{\omega_{n}+\omega_{n-1}}\,u(n)$. Then
$({\mathcal{H}}_{\omega,N}-\tilde{E})w=-\alpha_{N-1}w(N)e_{N-1}\implies\sqrt{\frac{\omega_{0}+\omega_{-1}}{\omega_{N}+\omega_{N-1}}}=-\alpha_{N-1}u(N)G^{E}_{\omega,N}(0,N-1).$
@@
$u(N)=\sqrt{\frac{\omega_{0}+\omega_{-1}}{\omega_{N}+\omega_{N-1}}}\det(\tilde{E}-{\mathcal{H}}_{\omega,N})\prod_{i=0}^{N-1}\alpha_{i}^{-1}.$
Let now $v$ be the solution satisfying the intitial conditions v(1)=1v(-1)=1 and $v(0)=0$. Then $\omega_{0}v(1)/\omega_{-1}=1$. Therefore,
$\frac{\omega_{0}}{\omega_{-1}}v(N)=\sqrt{\frac{\omega_{1}+\omega_{0}}{\omega_{N}+\omega_{N-1}}}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-1})\prod_{i=1}^{N-1}\alpha_{i}^{-1}.$
@@
$A^{E}_{N}(\omega)=\sqrt{\frac{\omega_{N-1}}{\omega_{-1}}}\begin{pmatrix}\frac{1}{\sqrt{\omega_{N}+\omega_{N-1}}}&0\\ &\\ 0&\frac{1}{\sqrt{\omega_{N-1}+\omega_{N-2}}}\end{pmatrix}\,\tilde{A}^{E}_{N}(\omega)\begin{pmatrix}\sqrt{\omega_{0}+\omega_{-1}}&0\\ &\\ 0&\sqrt{\omega_{-1}+\omega_{-2}}\end{pmatrix}$, (3.15)
where
$\tilde{A}^{E}_{N}(\omega)=\begin{pmatrix}\det(\tilde{E}-{\mathcal{H}}_{\omega,N})\prod_{i=0}^{N-1}\alpha_{i}^{-1}&-\alpha_{-1}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-1})\prod_{i=0}^{N-1}\alpha_{i}^{-1}\\ \,&\,\\ \det(\tilde{E}-{\mathcal{H}}_{\omega,N-1})\prod_{i=0}^{N-2}\alpha_{i}^{-1}&-\alpha_{-1}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-2})\prod_{i=0}^{N-2}\alpha_{i}^{-1}\end{pmatrix}$.
@@
Assume that $\mu$ has a bounded distortion property. Then there is a constant $C\geqslant 1$ such that
@@
Proof of Proposition 3.7. The proof follows a similar logic to that in [2], with the modification that we occasionally replace the operator $H_{\omega}$ by ${\mathcal{H}}_{\omega}$ and the point $E$ by $\tilde{E}=2\cos\sqrt{E}$.
Let $N$ and $\varepsilon$ be given. Fix $s\in{\mathbb{Z}}$ and $K\geqslant\max\{N,\ln^{2}(|s|+1)\}$. Define the set $D_{K,s}$ as the set of $\omega$ for which (3.24),(3.25) are satisfied for some $N_{1}$, $N_{2}$, $r$, and $m$ obeying the conditions $0\leqslant N_{1},N_{2}\leqslant K^{9}$, $K^{10}\leqslant r\leqslant\bar{K}$, and $m\in\{K,2K\}$. Note that
@@
Proof. Again, we define $w(n)=\sqrt{\omega_{n}+\omega_{n-1}}u(n)$. Then for any interval $[a,b]$,
$w(n)=-G_{\omega,[a,b]}^{E}(n,a)\alpha_{a-1}w(a-1)-G^{E}_{\omega,[a,b]}(n,b)\alpha_{b}w(b+1)$ (3.44)
@@
$\begin{split}|w(\ell_{i}+s_{0})|\leqslant 2(|w(a_{i}+s_{0})|+|w(b_{i}+s_{0}-1)|)\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3})\\ \leqslant C_{u}(5+|a_{i}|+|b_{i}|)\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3})\\ \leqslant C_{u}(5+2K^{9})\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3}).\end{split}$
@@
Now we use the condition $|u(s_{0})|=1$ and apply (3.44) with $a=s_{0}+\ell_{1}+1$ and $b=s_{0}+\ell_{2}-1$ to get the estimate
$\begin{split}1\leqslant|w(s_{0})|\leqslant 2|G^{E}_{\omega,[a,b]}(s_{0},a)|\,|w(s_{0}+\ell_{1})|+2|G^{E}_{\omega,[a,b]}(s_{0},b)|\,|w(s_{0}+\ell_{2})|\\ \leqslant 2\bigl(|G^{E}_{\omega,[a,b]}(s_{0},a)|+|G^{E}_{\omega,[a,b]}(s_{0},b)|\bigr)e^{-2K^{2}}\leqslant 4\|G^{E}_{\omega,[a,b]}\|e^{-2K^{2}}.\end{split}$
@@
To proceed, we introduce the following lemma, known as the Avalanche Principle, developed by Goldstein and Schlag to study the Lyapunov exponents of Schrödinger cocycles.
@@
for all sufficiently large $N_{0}$. This implies the relation